# Supermassive black holes swallow stellar objects at high rates: from Little Red Dots to Black Hole Stars

### v2

@@ Abstract
Supermassive black hole growth plausibly occurs via runaway astrophysical black hole mergers in nuclear star clusters that form intermediate mass black hole seeds at high redshifts. Such a model of Little Red Dots yields an order-of-magnitude higher rate of tidal disruption events than that of black hole captures. Our prediction, normalised to our proposed resolution of SMBH seeding, yields detectable TDE rates at high redshift. The resulting dense gas cocoons generate the nuclei of LRDs, each incorporating a central massive black-hole-star, with comparable masses in gas, stars, and massive black hole within a scale of around a parsec as inferred from the various spectral signatures.
@@ I Introduction
Extreme nuclear transients (ENTs) are rare, ultra-luminous flares ($10^{45}\,\rm erg\,s^{-1}$), exceeding supernova energies and more easily detected at $z>1$ than classical tidal disruption events (TDEs) in flux-limited surveys. The JWST has also uncovered numerous “Little Red Dots” (LRDs; $4<z<8$), likely AGN powered by $\sim 10^{7}$$10^{8}\,M_{\odot}$ supermassive black holes (SMBHs) (Matthee and others, 2024; Maiolino and others, 2024; Greene and others, 2024).
@@
Typical TDE durations are $3\times 10^{6}$$10^{7}\,\rm sec$ (Melchor et al., 2025). Rates are $\sim 10^{-5}\,\rm galaxy^{-1}\,yr^{-1}$ in normal galaxies, enhanced by 10–100 times in post-starburst systems (van Velzen et al., 2020). A new class of ambiguous nuclear transients (ANTs) is also observed (Wiseman and others, 2025).
Bellovary (2025) estimate a high-$z$ TDE rate of $\sim 10^{-4}\,\rm yr^{-1}$ (for $5<z<8$) by linking LRDs to SMBH-seed densities assuming purely TDE-powered emission, a rate too low for SMBH growth and implying LRD masses below earlier claims. A $\sim 10^{7}\,M_{\odot}$ BH has meanwhile been directly measured in a lensed $z=7$ LRD (Juodžbalis et al., 2025). Since LRDs are likely AGN-dominated, Kritos and Silk (2025) show that including SMBH-growth needs (mergers + accretion) significantly boosts high-$z$ TDE rates, which we argue here may be observable.
In this work, we consider an N-body snapshot of an NSC with a central SMBH and compute the instantaneous TDE and compact-object capture rates in the system, accounting for loss-cone physics and stellar evolution.
@@ II Stellar feeding onto central hole
We compute event rates with Monte Carlo simulations in the Keplerian potential $\Phi=-GM_{\rm SMBH}/r$ within the influence radius $r_{\rm infl}$, containing $N_{\star}\simeq 2M_{\rm SMBH}/\overline{m}_{\star}$ stars. Stellar masses are drawn from an initial mass function (IMF) and evolved with updated-BSE (Banerjee et al., 2020). Assuming $n_{\star}\propto r^{-\gamma}$ we sample radii from $p(r)\propto r^{3-\gamma}$ and orbits from the isotropic distribution $f(a,e)\propto a^{2-\gamma}e$ (Merritt, 2013, Eq. 4.36). Orbital periods follow from Kepler’s law, $P\simeq 2\pi GM_{\rm SMBH}m_{\star}(-E)^{-3/2}$.
Angular-momentum diffusion is driven by non-resonant relaxation (NRR) with a timescale (Merritt, 2013, Eq. 5.61)
@@
where $\sigma_{\star}=(1+\gamma)^{-1/2}(GM_{\rm SMBH}/r)^{1/2}$ and $\ln\Lambda\simeq\ln N_{\star}$. In the Keplerian regime, we also include resonant relaxation (RR) (Rauch and Tremaine, 1996; Gurkan and Hopman, 2007), with an effective relaxation time $t_{\rm R}=(t_{\rm NRR}^{-1}+t_{\rm RR}^{-1})^{-1}$ (Hopman and Alexander, 2006a, Eq. 11).
Feeding rates are computed using the loss-cone theory by Syer and Ulmer (1999). Defining $\theta_{\rm D}=(P/t_{\rm R})^{1/2}$ (Frank and Rees, 1976, Eq. 12), the loss cone is empty for $\theta_{\rm lc}>\theta_{\rm D}$, with rate $\ln(2/\theta_{\rm lc})t_{\rm R}]^{-1}$, and full otherwise, with rate $\theta_{\rm lc}^{2}/P$ per unit star. Summing over stars within $r_{\rm infl}$ yields the total instantaneous rate $\Gamma_{\rm lc}(t|M_{\rm SMBH})$.
For compact objects, we additionally include capture by gravitational bremsstrahlung (Quinlan and Shapiro, 1989, Eq. 11) and GW-driven inspirals (EMRIs). The effective loss-cone radius is the maximum of tidal, bremsstrahlung, and marginally bound radii. For EMRIs, $\theta_{\rm D}=(t_{\rm GW}/t_{\rm R})^{1/2}$ and the full-loss-cone regime is absent (Alexander and Hopman, 2003). We compute $t_{\rm GW}$ using the fit of Mandel (2021).
@@ III Initial conditions
Five representative time epochs ($10^{0}$, $10^{1}$, $10^{2}$, $10^{3}$, $10^{4}\,\rm Myr$) and three SMBH masses ($10^{5},\,10^{6},\,10^{7}\,M_{\odot}$) are considered with $a_{\rm SMBH}=0.5$. We link $r_{\rm infl}=GM_{\rm SMBH}/\sigma^{2}$ to $M_{\rm SMBH}$ through the $M_{\rm SMBH}$$\sigma$ relation (McConnell et al., 2011, Fig. 3). We assume the SMBH influence region is fully populated ($M_{\star}\gtrsim M_{\rm SMBH}$); otherwise, rates would scale down accordingly.
@@ IV Mass accretion rate
Figure 1: Temporal mass accretion rates from captures onto $10^{5}\,M_{\odot}$ (dotted), $10^{6}\,M_{\odot}$ (solid), and $10^{7}\,M_{\odot}$ (dashed) SMBHs, decomposed by stellar type via loss-cone contributions. Abbreviations: MS (main sequence); WD (white dwarf); NS (neutron star); BH (black hole). Points with error bars indicate mean values and standard deviations from ten realizations. The inset displays the duty cycle of MS (red) and giant (blue) stars.
Figure 1 shows the SMBH mass accretion rate by stellar type. Means and standard deviations are computed from ten realizations per time bin. Assuming a fraction $f_{\rm acc}$ of a stellar mass $m$ is accreted upon loss-cone entry, the accretion rate is $\dot{M}_{\rm SMBH}=f_{\rm acc}\,m\,\Gamma_{\rm lc}^{\rm(A)}(t|M_{\rm SMBH})$ where $\rm A$ is the stellar class. We adopt $f_{\rm acc}=0.5$ for MS stars (TDEs), $f_{\rm acc}=0.5$ for giants (envelope disruption), and $f_{\rm acc}=1$ for compact remnants (BHs, NSs, WDs).
@@
Integrating $\dot{M}_{\rm SMBH}$ over $t\in[0,1]\,\rm Gyr$ yields total mass growths of $\simeq 3.9$, $0.45$, and $0.057$ times the initial SMBH mass for $10^{5}$, $10^{6}$, and $10^{7}\,M_{\odot}$, respectively. The cumulative growth scales approximately as $\propto\sqrt{t}$ (also Stone et al., 2017). These values represent upper limits, as relaxation-driven ejections are neglected.
@@ V Flaring events
Figure 2: Peak luminosity versus characteristic rest-frame time-scale of astronomical transients, with peak absolute magnitudes normalized to a solar V-band magnitude of 4.83; ${\rm M}_{\rm pk}=2.5\log_{10}(L_{\rm pk}/L_{\odot})-4.83$. Tidal disruption flares of MS stars by SMBHs of $10^{5}\,M_{\odot}$ (crosses), $10^{6}\,M_{\odot}$ (circles), and $10^{7}\,M_{\odot}$ (diamonds) are shown as colored symbols, with the color bar indicating stellar mass. Five distinct sets correspond to five evolutionary times (bold labels). The six red symbols denote observed flares with $L_{\rm pk}>10^{45}\,\rm erg\,s^{-1}$. Gray dotted horizontal lines mark the Eddington luminosities for the labeled SMBH masses. Abbreviations: ENTs (extreme nuclear transients); SLSNe (superluminous supernovae); ANTs (ambiguous nuclear transients); TDEs (tidal disruption events); ccSNe (core-collapse supernovae); IaSNe (Type Ia supernovae) following the Phillips relationship.
The SMBH disrupts MS stars, while giant-star TDEs are typically partial and involve only envelope stripping (Navarro and Piran, 2025). The characteristic flare timescale is the fallback time (Gezari, 2021, Eq. 4),
@@
corresponding to the orbital period of the most bound debris. We adopt the shock-powered emission model of Krolik et al. (2025), in which stream self-intersections near apocenter produce radiation with efficiency (Krolik et al., 2025, Eq. 10)
@@
In the efficient-cooling limit ($t_{\rm cool}\ll t_{0}$), the peak luminosity equals the dissipation rate (Krolik et al., 2025, Eq. 16),
@@
well below $L_{\rm Edd}\simeq 1.2\times 10^{45}\,{\rm erg\,s^{-1}}(M_{\rm SMBH}/10^{6}\,M_{\odot})$. For inefficient cooling ($t_{\rm cool}\gg t_{0}$), $L_{\rm pk}\sim L_{\rm Edd}$ and the flare duration is $t_{0}\simeq m_{\star}c^{2}/L_{\rm pk}$ (Krolik et al., 2025, Eq. 18).
Figure 2 shows the peak luminosities and rest-frame decay times of MS TDEs in the efficient-cooling limit, colored by stellar mass, alongside observed transient classes and six extreme flares (Hinkle et al., 2025; Graham and others, 2025). The most luminous events arise from the most massive MS stars. As the stellar population evolves, the maximum MS mass decreases while stellar radii increase near turnoff, producing a late-time turnover since $t_{0}\propto R_{\star}^{3/2}$. TDEs of $\sim 30$$150\,M_{\odot}$ MS stars by $10^{6}$$10^{7}\,M_{\odot}$ SMBHs are consistent with ENTs, while lower-mass systems fall in the ANT regime.
@@ VI Integration over cosmological volume
We focus on the redshift range $z<6$. Assuming a constant SMBH mass function with comoving number density $n_{\rm SMBH}$, we weight three representative SMBH masses by $w_{j}\propto M_{{\rm SMBH},j}^{-2}$ and normalize $\sum_{j}w_{j}=1$. The source-frame loss-cone rate density for class A is
@@
including both captures and inspirals, and assuming a uniform time prior.
The observer-frame rate follows from integrating over comoving volume and applying cosmological time dilation (Kritos and Silk, 2025, Eq. 27). Assuming ${\cal R}_{\rm lc}^{(\rm A)}$ is constant over $0\leq z\leq 6$,
@@ A ${\cal R}_{\rm capture}$ $R_{\rm capture}$ ${\cal R}_{\rm inspiral}$ $R_{\rm inspiral}$
which is on the same order of Sołtan’s estimated mass density ($8\times 10^{13}\,M_{\odot}\,\rm Gpc^{-3}$, Soltan, 1982), implying gas accretion dominates SMBH growth but loss-cone effects contribute at the $\sim 25\%$ level.
We find an MS TDE rate of $\approx 5\times 10^{3}\,{\rm Gpc^{-3}\,yr^{-1}}$ at $4\leq z\leq 6$, with only a small detectable fraction (e.g. UVEX; Kulkarni and others 2021) while inspirals are negligible in all classes. For comparison, the local TDE rate density is $\sim 15$$200\,{\rm Gpc^{-3}\,yr^{-1}}$ (van Velzen and Farrar, 2014), below theoretical $z\sim 0$ expectations (Stone and Metzger, 2016), while recent candidates at $z>1$ (Gu et al., 2025) imply rates of $\sim 20$$80\,{\rm Gpc^{-3}\,yr^{-1}}$ at $z\sim 1$ (Graham and others, 2025). In addition, ${\cal R}_{\rm MS}:{\cal R}_{\rm giant}\sim 30$. Finally, stellar remnants capture at intrinsic rates of hundreds per year, and about two BH-EMRI every century, producing GW sources, which we discuss next.
@@ VII Gravitational-wave background
We compute the characteristic GW strain of each capture, $h_{c}(f)$, at the observer-frame frequency $f$, using Eq. (17) of Bonetti and Sesana (2020), including the first 100 harmonics. Merger–ringdown and spin effects are neglected. Example individual sources at $z\sim 5$ are shown in the inset of Fig. 3 together with the sensitivities of LGWA (Harms and others, 2021) and LISA (Amaro-Seoane and others, 2017). Captures originate from highly eccentric ($e\simeq 1^{-}$) orbits that rapidly circularize and merge within $\simeq 28\,{\rm min}\,M_{\rm SMBH}/(10^{6}\,M_{\odot})({10\,\rm km\,s^{-1}}/\sigma)^{3}$ (O’Leary et al., 2009). We find that $\simeq 54\%$ ($\simeq 1.3\%$), $\simeq 7.9\%$ ($\simeq 0.6\%$), and $\simeq 3.3\%$ ($\simeq 0.5\%$) of all BH, NS, and WD captures, respectively, have $\rm SNR>10$ in LISA (LGWA).
The cumulative contribution of captures produces a stochastic GW background (SGWB), whose dimensional energy density over an observation time $T_{\rm obs}$ we approximate as (Belgacem et al., 2025, Eq. 3.83)
@@
The resulting SGWBs are shown in Fig. 3 (thick lines). The WD capture background peaks at deciHz and are observable with an $\rm SNR>10$ when compared with the corresponding LGWA PLS (Ajith and others, 2025) and LISA PLS (Schmitz, 2020) for $T_{\rm obs}=4\,\rm yr$. For comparison, we also show the background from massive black-hole mergers under two seeding scenarios (Çalışkan et al., 2025). While the BH and NS backgrounds have $\rm SNR<10$ within four years, there are still thousands of captures within $z<6$ with LISA. Finally, the SGWB from inspirals is undetectable.
VIII Formation of the black-hole star
We adopt a BH–star framework in which the SMBH is embedded in a dense cocoon of gas and stars. This model provides a consistent interpretation of LRDs, satisfies spectral constraints, and enables rapid SMBH growth via bursty, super-Eddington gas accretion (Kritos and Silk, 2025; Kokorev et al., 2025). The buildup of dense gas cocoons produces extreme Balmer breaks, Balmer absorption, and $H\beta$ emission, identifying LRDs as BH-stars (de Graaff et al., 2025; Naidu and others, 2025); see also (Begelman and Dexter, 2026; Nandal and Loeb, 2025; Rusakov et al., 2025).

### v3

# Toward Black Hole Stars: supermassive black hole growth in nuclear clusters via stellar-object and gas accretion

@@ Abstract
Supermassive black hole (SMBH) growth plausibly occurs via runaway astrophysical black hole mergers in nuclear star clusters that form intermediate mass black hole seeds at high redshifts. Such a model yields an order-of-magnitude higher rate of tidal disruption events than that of compact-object captures. Our prediction, normalized to our proposed resolution of SMBH seeding, yields detectable tidal disruption event rates at high redshift. The resulting dense gas cocoons generate compact galactic nuclei, each incorporating a central, massive, black hole star, with comparable masses in gas, stars, and massive black holes within a scale of around a parsec as inferred from the various Little Red Dot spectral signatures.
@@ I Introduction
Extreme nuclear transients (ENTs) are rare, ultraluminous flares ($10^{45}\,\rm erg\,s^{-1}$), exceeding supernova energies and more easily detected at $z>1$ than classical tidal disruption events (TDEs) in flux-limited surveys. The JWST has also uncovered numerous “Little Red Dots” (LRDs; $4<z<8$), likely AGN powered by $\sim 10^{7}$$10^{8}\,M_{\odot}$ supermassive black holes (SMBHs; Greene and others, 2024; Maiolino and others, 2024; Matthee and others, 2024).
@@
Typical TDE durations are $3\times 10^{6}$$10^{7}\,\rm sec$ (Melchor et al., 2025). Rates are $\sim 10^{-5}\,\rm galaxy^{-1}\,yr^{-1}$ in normal galaxies, enhanced by 10–100 times in poststarburst systems (van Velzen et al., 2020). A new class of ambiguous nuclear transients (ANTs) is also observed (Wiseman and others, 2025).
Bellovary (2025) estimates a high-$z$ TDE rate of $\sim 10^{-4}\,\rm yr^{-1}$ (for $5<z<8$) by linking LRDs to SMBH-seed densities assuming purely TDE-powered emission, a rate too low for SMBH growth and implying LRD masses below earlier claims. A $\sim 10^{7}\,M_{\odot}$ black hole (BH) has meanwhile been directly measured in a lensed $z=7$ LRD (Juodžbalis et al., 2025). Since LRDs are likely AGN-dominated, Kritos and Silk (2025) show that including SMBH-growth needs (mergers + accretion) significantly boosts high-$z$ TDE rates, which we argue here may be observable.
In this work, we analyze an $N$-body snapshot at time $t$ of a star cluster within a central SMBH’s sphere of influence by sampling its spherically symmetric distribution function without time evolution. From these snapshots, we compute the instantaneous TDE and compact-object capture rates, incorporating loss-cone physics and stellar evolution.
@@ II Stellar feeding onto central hole
We compute event rates with Monte Carlo simulations in the Keplerian potential $\Phi=-GM_{\rm SMBH}/r$ within the influence radius $r_{\rm infl}$, containing $N_{\star}\simeq 2M_{\rm SMBH}/\overline{m}_{\star}$ stars. Stellar masses are drawn from an initial mass function (IMF) and evolved with updated-BSE (Banerjee et al., 2020). Assuming $n_{\star}\propto r^{-\gamma}$ we sample radii from $p(r)\propto r^{3-\gamma}$ and orbits from the isotropic distribution $f(a,e)\propto a^{2-\gamma}e$ (Merritt, 2013, Eq. (4.36)). Orbital periods follow from Kepler’s law, $P\simeq 2\pi GM_{\rm SMBH}m_{\star}(-E)^{-3/2}$.
Angular-momentum diffusion is driven by nonresonant relaxation (NRR) with a timescale (Merritt, 2013, Eq. (5.61))
@@
where $\sigma_{\star}=(1+\gamma)^{-1/2}(GM_{\rm SMBH}/r)^{1/2}$ and $\ln\Lambda\simeq\ln N_{\star}$. In the Keplerian regime, we also include resonant relaxation (RR; Rauch and Tremaine, 1996; Gurkan and Hopman, 2007), with an effective relaxation time $t_{\rm R}=(t_{\rm NRR}^{-1}+t_{\rm RR}^{-1})^{-1}$ (Hopman and Alexander, 2006a, Eq. 11).
Feeding rates are computed using the loss-cone theory by Syer and Ulmer (1999). Defining $\theta_{\rm D}=(P/t_{\rm R})^{1/2}$ (Frank and Rees, 1976, Eq. (12)), the loss cone is empty for $\theta_{\rm lc}>\theta_{\rm D}$, with rate $\ln(2/\theta_{\rm lc})t_{\rm R}]^{-1}$, and full otherwise, with rate $\theta_{\rm lc}^{2}/P$ per unit star. Summing over stars within $r_{\rm infl}$ yields the total instantaneous rate $\Gamma_{\rm lc}(t|M_{\rm SMBH})$.
For compact objects, we additionally include capture by gravitational bremsstrahlung (Quinlan and Shapiro, 1989, Eq. (11)) and gravitational wave (GW)-driven inspirals, known as extreme mass-ratio inspirals (EMRIs). The effective loss-cone radius is the maximum of tidal, bremsstrahlung, and marginally bound radii. For EMRIs, $\theta_{\rm D}=(t_{\rm GW}/t_{\rm R})^{1/2}$ and the full-loss-cone regime is absent (Alexander and Hopman, 2003). We compute $t_{\rm GW}$ using the fit of Mandel (2021).
@@ III Initial conditions
Five representative time epochs ($10^{0}$, $10^{1}$, $10^{2}$, $10^{3}$, and $10^{4}\,\rm Myr$) and three SMBH masses ($10^{5},\,10^{6},\ {\rm and}\ 10^{7}\,M_{\odot}$) are considered with $a_{\rm SMBH}=0.5$. We link $r_{\rm infl}=GM_{\rm SMBH}/\sigma^{2}$ to $M_{\rm SMBH}$ through the $M_{\rm SMBH}$$\sigma$ relation (McConnell et al., 2011, Fig. 3). We assume the SMBH influence region is fully populated ($M_{\star}\gtrsim M_{\rm SMBH}$); otherwise, rates would scale down accordingly.
@@ IV Mass accretion rate
Figure 1: Temporal mass accretion rates from captures onto $10^{5}$ (dotted), $10^{6}$ (solid), and $10^{7}\,M_{\odot}$ (dashed) SMBHs, decomposed by stellar type via loss-cone contributions. Abbreviations: MS (main sequence); WD (white dwarf); NS (neutron star); BH (black hole). Points with error bars indicate mean values and standard deviations from 10 realizations. The inset displays the duty cycle of MS (red) and giant (blue) stars.
Figure 1 shows the SMBH mass accretion rate by stellar type. Means and standard deviations are computed from 10 realizations per time bin. Assuming a fraction $f_{\rm acc}$ of a stellar mass $m$ is accreted upon loss-cone entry, the mass accretion rate per unit star is $d\dot{M}_{\rm SMBH}^{(\rm A)}/dN_{\rm A}=f_{\rm acc}\,m\,d\Gamma_{\rm lc}^{\rm(A)}(t|M_{\rm SMBH})/dN_{\rm A}$ where $\rm A$ is the stellar class. We adopt $f_{\rm acc}=0.5$ for MS stars (TDEs), $f_{\rm acc}=0.5$ for giants (envelope disruption), and $f_{\rm acc}=1$ for compact remnants (BHs, NSs, WDs).
@@
Integrating $\dot{M}_{\rm SMBH}$ over $t\in[0,1]\,\rm Gyr$ yields total mass growths of $\simeq 3.9$, $0.45$, and $0.057$ times the initial SMBH mass for $10^{5}$, $10^{6}$, and $10^{7}\,M_{\odot}$, respectively. The cumulative growth scales approximately as $\propto\sqrt{t}$ (also see Stone et al., 2017). These values represent upper limits, as relaxation-driven ejections are neglected.
@@ V Flaring events
Figure 2: Peak luminosity versus characteristic rest-frame time-scale of astronomical transients, with peak absolute magnitudes normalized to a solar $V$-band magnitude of 4.83; ${\rm M}_{\rm pk}=2.5\log_{10}(L_{\rm pk}/L_{\odot})-4.83$. Tidal disruption flares of MS stars by SMBHs of $10^{5}$ (crosses), $10^{6}$ (circles), and $10^{7}\,M_{\odot}$ (diamonds) are shown as colored symbols, with the color bar indicating stellar mass. Five distinct sets correspond to five evolutionary times (bold labels). The six red symbols denote observed flares with $L_{\rm pk}>10^{45}\,\rm erg\,s^{-1}$. Gray dotted horizontal lines mark the Eddington luminosities for the labeled SMBH masses. Abbreviations: ENTs (extreme nuclear transients); SLSNe (superluminous supernovae); ANTs (ambiguous nuclear transients); TDEs (tidal disruption events); ccSNe (core-collapse supernovae); Ia SNe (Type Ia supernovae) following the Phillips relationship.
The SMBH disrupts MS stars, while giant-star TDEs are typically partial and involve only envelope stripping (Navarro and Piran, 2025). The characteristic flare timescale is the fallback time (Gezari, 2021, Eq. (4)),
@@
corresponding to the orbital period of the most bound debris. We adopt the shock-powered emission model of Krolik et al. (2025), in which stream self-intersections near the apocenter produce radiation with efficiency (Krolik et al., 2025, Eq. (10))
@@
In the efficient-cooling limit ($t_{\rm cool}\ll t_{0}$), the peak luminosity equals the dissipation rate (Krolik et al., 2025, Eq. (16)),
@@
well below $L_{\rm Edd}\simeq 1.2\times 10^{45}\,{\rm erg\,s^{-1}}(M_{\rm SMBH}/10^{6}\,M_{\odot})$. For inefficient cooling ($t_{\rm cool}\gg t_{0}$), $L_{\rm pk}\sim L_{\rm Edd}$ and the flare duration is $t_{0}\simeq m_{\star}c^{2}/L_{\rm pk}$ (Krolik et al., 2025, Eq. (18)).
Figure 2 shows the peak luminosities and rest-frame decay times of MS TDEs in the efficient-cooling limit, colored by stellar mass, alongside observed transient classes and six extreme flares (Graham and others, 2025; Hinkle et al., 2025). The most luminous events arise from the most massive MS stars. As the stellar population evolves, the maximum MS mass decreases while stellar radii increase near turnoff, producing a late-time turnover since $t_{0}\propto R_{\star}^{3/2}$. TDEs of $\sim 30$$150\,M_{\odot}$ MS stars by $10^{6}$$10^{7}\,M_{\odot}$ SMBHs are consistent with ENTs, while lower-mass systems fall in the ANT regime.
@@ VI Integration over cosmological volume
We focus on the redshift range $z<6$. Assuming a constant SMBH mass function with comoving number density $n_{\rm SMBH}$, we weight three representative SMBH masses by $w_{j}\propto M_{{\rm SMBH},j}^{-2}$ and normalize $\sum_{j}w_{j}=1$. Our choice of equal SMBH mass per unit logarithmic SMBH mass bin in the examined mass range is motivated by Gallo and Sesana (2019) and Greene et al. (2020). The source-frame loss-cone rate density for class A is
@@
including both captures and inspirals, and assuming a uniform time prior over which index $k$ runs.
The observer-frame rate follows from integrating over comoving volume and applying cosmological time dilation (Kritos and Silk, 2025, Eq. (27)). Assuming ${\cal R}_{\rm lc}^{(\rm A)}$ is constant over $0\leq z\leq 6$,
@@ A ${\cal R}_{\rm capture}$ $R_{\rm capture}$ ${\cal R}_{\rm inspiral}$ $R_{\rm inspiral}$
which is on the same order as Sołtan’s estimated mass density ($8\times 10^{13}\,M_{\odot}\,\rm Gpc^{-3}$; Soltan, 1982), implying gas accretion dominates SMBH growth but loss-cone effects contribute at the $\sim 25\%$ level.
We find an MS TDE rate of $\approx 5\times 10^{3}\,{\rm Gpc^{-3}\,yr^{-1}}$ at $4\leq z\leq 6$, with only a small detectable fraction (e.g., UVEX; Kulkarni and others 2021) while inspirals are negligible in all classes. For comparison, the local TDE rate density is $\sim 15$$200\,{\rm Gpc^{-3}\,yr^{-1}}$ (van Velzen and Farrar, 2014), below theoretical $z\sim 0$ expectations (Stone and Metzger, 2016), while recent candidates at $z>1$ (Gu et al., 2025) imply rates of $\sim 20$$80\,{\rm Gpc^{-3}\,yr^{-1}}$ at $z\sim 1$ (Graham and others, 2025). In addition, ${\cal R}_{\rm MS}:{\cal R}_{\rm giant}\sim 30$. Finally, stellar remnants capture at intrinsic rates of hundreds per year, and about two BH-EMRIs every century, producing GW sources, which we discuss next.
@@ VII Gravitational-wave background
We compute the characteristic GW strain of each capture, $h_{c}(f)$, at the observer-frame frequency $f$, using Eq. (17) of Bonetti and Sesana (2020), including the first 100 harmonics. Merger–ringdown and spin effects are neglected. Example individual sources at $z\sim 5$ are shown in the inset of Fig. 3 together with the sensitivities of LGWA (Harms and others, 2021) and LISA (Amaro-Seoane and others, 2017). Captures originate from highly eccentric ($e\simeq 1^{-}$) orbits that rapidly circularize and merge within $\simeq 28\,{\rm min}\,M_{\rm SMBH}/(10^{6}\,M_{\odot})({10\,\rm km\,s^{-1}}/\sigma)^{3}$ (O’Leary et al., 2009). We find that $\simeq 54\%$ ($\simeq 1.3\%$), $\simeq 7.9\%$ ($\simeq 0.6\%$), and $\simeq 3.3\%$ ($\simeq 0.5\%$) of all BH, NS, and WD captures, respectively, have signal-to-noise ratio (SNR) >10>10 in LISA (LGWA).
The cumulative contribution of captures produces a stochastic GW background (SGWB), whose dimensional energy density over an observation time $T_{\rm obs}$ we approximate as (Belgacem et al., 2025, Eq. (3.83))
@@
The resulting SGWBs are shown in Fig. 3 (thick lines). The WD capture background peaks at dHz and are observable with an $\rm SNR>10$ when compared with the corresponding LGWA Power-Law integrated Sensitivity (PLS; Ajith and others, 2025) and LISA PLS (Schmitz, 2020) for $T_{\rm obs}=4\,\rm yr$. For comparison, we also show the background from massive BH mergers under two seeding scenarios (Çalışkan et al., 2025). While the BH and NS backgrounds have $\rm SNR<10$ within $\rm 4\,yr$, there are still thousands of captures within $z<6$ with LISA. Finally, the SGWB from inspirals is undetectable.
VIII Formation of the black hole star
[…]